Now, It's Your Turn 1. Create your own or similar real-life situation where rational function is applied 2. In a bond paper, present the problem from Field of application up to Presentation as illustrated above. 3. You can use graphing paper, MS Excel, Desmos, or any graphing app to graph the function. ​


1
Expert's answer
2021-11-15T17:37:40-0500

from real-life situation:

equation of distance travelled as function of time:

"x(t)=at^2\/2+v_0t"

where a is acceleration,

"v_0" is the initial speed.


for example, if a car has

"a=1" m/s2 and "v_0=20" m/s

then

"x(t)=t^2\/2+20t"